# Picosecond Wireless Synchronization with Entangled Photons via Grid-Based Quantum Coverage in Indoor Optical Systems

### v1

@@ V Simulations
To assess the synchronization performance under realistic quantum indoor environments, we conducted Monte-Carlo simulations based on the system model described in Sections IIIV. The room was configured as a rectangular indoor space with dimensions $L=6\,\text{m}$, $W=6\,\text{m}$, and $H=3\,\text{m}$. The quantum optical coverage was set at height $h_{\text{cov}}=1\,\text{m}$ above the floor, and the grid-based beam steering was evaluated for multiple spatial configurations with $N_{g}\in\{25,100,225\}$, corresponding to $5\times 5$, $10\times 10$, and $15\times 15$ grid layouts. Each grid region was targeted by a distinct optical beam, and photon propagation followed the diffraction and spatial uncertainty models described earlier.
@@
The synchronization sequence spanned a total duration $T_{\text{seq}}$ ranging from $1\,\mu\text{s}$ to $2\,\text{ms}$, divided into slots of duration $T_{qs}=10\,\text{ns}$. This results in a total number of time slots $N_{s}=T_{\text{seq}}/T_{qs}$, which varied between $10^{2}$ and $2\times 10^{5}$ depending on the experiment. Each slot could contain a single entangled photon pair with mean generation rate $\mu_{t}=0.5$, while the SPAD detector timing jitter was modeled as Gaussian noise with standard deviation $\sigma_{d}=200\,\text{ps}$. The SPDC source temporal jitter was modeled with standard deviation $\sigma_{t}=50\,\text{ps}$ [35]. Background photon arrivals followed a Poisson distribution with mean $\mu_{b}=5\times 10^{-5}$. The quantum detection efficiency of the SPAD was set to $\eta_{d}=0.6$, and the user receiver was modeled with a circular aperture of radius $r_{a}=0.02\,\text{m}$. The simulation parameters are summarized in Table II, where variables marked with a range were evaluated parametrically, and exact values for each scenario are specified in the respective figure captions.
TABLE II: Simulation Parameters for Quantum Synchronization System
@@
Another critical parameter that significantly affects the performance of wireless quantum synchronization systems is the mean photon-pair generation rate per slot, denoted by $\mu_{t}$. Fig. 3 illustrates the average synchronization error as a function of $\mu_{t}$ over a wide range from $0.1$ to $4$, under fixed environmental and timing configurations. The simulation considers a $15\times 15$ grid structure ($N_{g}=225$) and a synchronization duration of $T_{\text{seq}}=1\,\text{ms}$ per user. All other system parameters follow the nominal values specified in Table II. As evident in the results, the value of $\mu_{t}$ plays a dual and non-trivial role in synchronization accuracy. When $\mu_{t}$ is too small, the probability that the SPDC source generates a valid single photon pair in a given quantum slot decreases sharply, leading to high rates of slot failure and consequently degraded synchronization performance due to insufficient valid timestamps. Conversely, if $\mu_{t}$ becomes too large, the probability of generating two or more photon pairs in the same slot increases, which, according to the protocol defined in (17), renders the slot invalid for synchronization purposes. Such multi-pair emissions introduce ambiguity and break the one-to-one correspondence between reference and user photons. Therefore, there exists an optimal range for $\mu_{t}$ that balances the trade-off between too few photon events and too many invalid multiphoton occurrences. As shown in Fig. 3, this balance is achieved near $\mu_{t}\approx 1$, which minimizes the average synchronization error. Selecting $\mu_{t}$ carefully around this optimal region is essential for reliable and high-precision quantum synchronization under constrained photon budgets and noisy channel conditions.
VI Conclusion and Future Work
In this work, we introduced a grid-structured, entanglement-based quantum synchronization architecture for indoor wireless optical systems. The proposed framework leverages a central ceiling-mounted SPDC transmitter that steers time-energy entangled photons toward spatially targeted beams, each corresponding to a predefined spatial grid. By estimating the user’s location and aligning the optical beam accordingly, the system increases the likelihood of photon reception despite the probabilistic nature of quantum propagation and detection. A beam-local coordinate transformation was applied to analytically model the spatial photon reception probability, which plays a key role in determining synchronization reliability. To overcome sparsity and randomness in photon arrivals, we proposed a robust two-stage synchronization recovery algorithm based on sparse bit-pattern matching and timestamp alignment. This method enables accurate estimation of clock offsets even in the presence of background noise, multiphoton events, and detector jitter. Simulation results demonstrated the effectiveness of the approach under practical conditions. It was shown that increasing the synchronization sequence duration significantly reduces both synchronization failure probability and mean timing error. Furthermore, increasing grid resolution from $5\times 5$ to $15\times 15$ notably improves photon alignment and enhances synchronization accuracy, achieving sub-100 ps mean error within $1\,\text{ms}$ of allocated time. Additionally, we evaluated the effect of photon-pair generation rate $\mu_{t}$ and identified an optimal range that minimizes timing error by balancing between insufficient single-pair events and excess multiphoton emissions.
Future work will focus on extending the system to multi-transmitter environments to enhance spatial coverage, support user mobility, and improve robustness against obstructions. Additionally, a promising direction involves integrating synchronization with quantum communication, sensing, and positioning into a unified framework. Such a system could enable sub-centimeter positioning and sub-picosecond synchronization accuracy, far surpassing the capabilities of conventional classical techniques, and pave the way for secure, high-precision, quantum-enabled indoor wireless infrastructures.

### v2

@@ VII Simulations
To assess the synchronization performance under realistic quantum indoor environments, we conducted Monte-Carlo simulations based on the system model described in Sections IIIV. The room was configured as a rectangular indoor space with dimensions $L=6\,\text{m}$, $W=6\,\text{m}$, and $H=3\,\text{m}$. The quantum optical coverage was set at height $h_{\text{cov}}=1\,\text{m}$ above the floor, and the grid-based beam steering was evaluated for multiple spatial configurations with $N_{g}\in\{25,100,225\}$, corresponding to $5\times 5$, $10\times 10$, and $15\times 15$ grid layouts [22]. Each grid region was targeted by a distinct optical beam, and photon propagation followed the diffraction and spatial uncertainty models described earlier.
@@
The synchronization sequence spanned a total duration $T_{\text{seq}}$ ranging from $1\,\mu\text{s}$ to $2\,\text{ms}$, divided into slots of duration $T_{qs}=10\,\text{ns}$. This results in a total number of time slots $N_{s}=T_{\text{seq}}/T_{qs}$, which varied between $10^{2}$ and $2\times 10^{5}$ depending on the experiment. Each slot could contain a single entangled photon pair with mean generation rate $\mu_{t}=0.5$, while the SPAD detector timing jitter was modeled as Gaussian noise with standard deviation $\sigma_{d}=200\,\text{ps}$. The SPDC source temporal jitter was modeled with standard deviation $\sigma_{t}=50\,\text{ps}$ [41]. Background photon arrivals followed a Poisson distribution with mean $\mu_{b}=5\times 10^{-6}$ [42]. The quantum detection efficiency of the SPAD was set to $\eta_{d}=0.6$, and the user receiver was modeled with a circular aperture of radius $r_{a}=0.02\,\text{m}$. Unless otherwise stated, the simulations assume an LOS channel and use the Gaussian detector jitter model. The simulation parameters are summarized in Table III, where variables marked with a range were evaluated parametrically, and exact values for each scenario are specified in the respective figure captions.
TABLE III: Simulation Parameters for Quantum Synchronization System
@@
Delay spred $\tau_{\text{rms}}$ $300\,\text{ps}$
VII-A Baseline Synchronization Performance under Ideal Conditions
@@
Another critical parameter that significantly affects the performance of wireless quantum synchronization systems is the mean photon-pair generation rate per slot, denoted by $\mu_{t}$. Fig. 3 illustrates the average synchronization error as a function of $\mu_{t}$ over a wide range from $0.1$ to $4$, under fixed environmental and timing configurations. The simulation considers a $15\times 15$ grid structure ($N_{g}=225$) and a synchronization duration of $T_{\text{seq}}=1\,\text{ms}$ per user. All other system parameters follow the nominal values specified in Table III. As evident in the results, the value of $\mu_{t}$ plays a dual and non-trivial role in synchronization accuracy. When $\mu_{t}$ is too small, the probability that the SPDC source generates a valid single photon pair in a given quantum slot decreases sharply, leading to high rates of slot failure and consequently degraded synchronization performance due to insufficient valid timestamps. Conversely, if $\mu_{t}$ becomes too large, the probability of generating two or more photon pairs in the same slot increases, which, according to the protocol defined in (17), renders the slot invalid for synchronization purposes. Such multi-pair emissions introduce ambiguity and break the one-to-one correspondence between reference and user photons. Therefore, there exists an optimal range for $\mu_{t}$ that balances the trade-off between too few photon events and too many invalid multiphoton occurrences. As shown in Fig. 3, this balance is achieved near $\mu_{t}\approx 1$, which minimizes the average synchronization error. Selecting $\mu_{t}$ carefully around this optimal region is essential for reliable and high-precision quantum synchronization under constrained photon budgets and noisy channel conditions.
@@
VII-B User Mobility, Handover, and Quasi-Static Assumption
User mobility is modeled under a quasi-static assumption, which is justified by the short duration of the synchronization window relative to typical indoor motion. For a synchronization interval $T_{\text{sync}}$ on the order of milliseconds (as evidenced in Fig. 2), and a maximum indoor walking speed $v\leq 1.5~\mathrm{m/s}$, the resulting displacement satisfies $\Delta d=vT_{\text{sync}}\ll w_{z}$, where $w_{z}$ denotes the beam footprint at the receiver plane. Consequently, user motion within a synchronization interval does not significantly affect the photon-reception probability or timing statistics. Mobility therefore manifests as a sequence of quasi-static intervals separated by grid handovers, rather than as continuous time variation. Upon handover, synchronization is re-established using the same procedure without reliance on prior state information, thereby validating the quasi-static assumption adopted in the simulations.
VII-C Impact of Spatial Uncertainty and Grid Resolution
In this subsection, we evaluate the robustness of the proposed synchronization framework under realistic user-positioning uncertainties. Positioning errors do not affect the photon generation or timing process itself; instead, their impact is purely geometric. Specifically, a mismatch between the estimated and true user location reduces the photon-reception probability by shifting the receiver away from the beam center, thereby decreasing the number of detected photon pairs available for synchronization.
To quantify this effect, four representative positioning error models are considered while maintaining the same RMS uncertainty $\sigma_{p}$ for fair comparison. Specifically, we consider: (i) a Gaussian model with independent zero-mean errors (discussed in Section III as the baseline), $n\sim\mathcal{N}(0,\sigma_{p}^{2}I)$; (ii) a Laplacian model with heavier tails, whose probability density is given by $f(x)=\frac{1}{2b}\exp(-|x|/b)$ with $b=\sigma_{p}/\sqrt{2}$; (iii) a spatially correlated Gaussian model reflecting correlation induced by anchor geometry in trilateration-based systems with covariance $\Sigma=\sigma_{p}^{2}\begin{bmatrix}1&\rho&0\\ \rho&1&0\\ 0&0&1\end{bmatrix}$; and (iv) a biased model combining Gaussian noise with a uniform offset, i.e., $n_{\text{Uniform}}\sim\mathcal{U}(-0.75\sigma_{p},\,0.75\sigma_{p})$, to emulate systematic localization errors. For each case, the photon-reception probability is computed using the beam-based reception model in (42), and the resulting synchronization error is obtained from the Cramér–Rao bound in (37).
Table  IV summarizes the resulting performance. The results show that different positioning error statistics primarily affect the photon-reception probability, while the timing statistics of individual detections remain unchanged. Even under the worst-case Laplacian model, the synchronization error increases by only 12%, requiring approximately 25% longer integration time to achieve the same precision. This demonstrates that the proposed synchronization method exhibits graceful degradation rather than abrupt failure. Importantly, the observed performance loss arises from reduced photon counts rather than increased timing noise. As a result, synchronization accuracy can be recovered by extending the observation window, confirming that the proposed framework is robust to realistic indoor localization errors and does not rely on idealized positioning assumptions.
TABLE IV: Synchronization performance under different user-positioning error models
Dist. $P_{\text{rec}}$ Sync Err. Time Penalty
Factor (dB)
Gaussian 0.0317 9.2 ps $1.00\times$ 0.0
Laplacian 0.0254 10.3 ps $1.25\times$ 1.0
Correlated 0.0280 9.8 ps $1.13\times$ 0.6
Biased 0.0299 9.5 ps $1.06\times$ 0.3
Simulation parameters: $w_{z}=0.283$ m, $r_{a}=0.02$ m, $\sigma_{p}=0.2$ m, $\sigma_{d}=200$ ps, $\eta_{d}=0.6$, $\rho$ = 0.7, and synchronization window of $1$ ms.
VII-D Impact of Multipath Dispersion and Non-Gaussian Timing Jitter
Fig. 4 illustrates the impact of multipath dispersion and non-Gaussian timing jitter on the synchronization accuracy as a function of the synchronization window length. Unlike the idealized LOS case, the synchronization error does not decay to sub 10 ps as the observation time increases. Instead, the error converges to a non-zero floor (i.e., RMS delay), which is a direct consequence of multipath-induced delay spread in the indoor channel. This behavior confirms that, beyond a certain integration duration, additional photon detections do not improve timing accuracy because the dominant impairment is no longer stochastic noise but a deterministic bias introduced by delayed reflected components. The effect is further exacerbated under Laplacian timing jitter, where the heavier tail slows convergence relative to the Gaussian case, even though the mean timing error remains bounded. Importantly, the observed error floor is not a limitation of the proposed synchronization algorithm but a physical consequence of the propagation environment. Since the multipath bias depends on room geometry, surface reflectivity, and transmitter–receiver placement, it can in principle be calibrated and compensated. After such calibration, the residual timing error remains in the picosecond regime, demonstrating that the proposed framework preserves high-precision synchronization performance even under realistic non-LOS and non-Gaussian conditions.
Figure 4: Average synchronization error as a function of the sequence duration $T_{\text{seq}}$ for several spatial grid sizes, evaluated under NLOS channel conditions and non‑Gaussian timing jitter.
VII-E Computational Complexity and Real-Time Feasibility
The computational implications of the proposed synchronization framework are also evaluated to assess its feasibility for real-time implementation. The two-stage synchronization algorithm is intentionally designed to be computationally efficient. The coarse alignment stage relies on sequence correlation, which can be implemented using FFT-based convolution with a computational complexity of $O(N_{s}\log N_{s})$, where $N_{s}$ denotes the number of time slots. For the largest sequence length considered in this work ($N_{s}=2\times 10^{5}$), this operation requires only a few milliseconds on contemporary DSP or FPGA platforms, consistent with reported real-time FFT performance [43]. The memory footprint is similarly modest, i.e., the binary detection sequence requires approximately $25\,\text{KB}$, while the associated timestamp vector occupies approximately $1.6\,\text{MB}$. Importantly, due to the sparse nature of photon arrivals, only a small fraction of slots contain valid detections, which significantly reduces the effective processing load in practice. As a result, the algorithm is well suited for real-time operation, including fast re-synchronization during user mobility and grid handovers. These results confirm that the proposed synchronization framework is computationally viable for practical indoor optical wireless deployments.
VIII Conclusion and Future Work
In this work, we presented a unified framework for entanglement assisted time synchronization in indoor optical wireless systems, with particular emphasis on grid based architectures that support mobility and spatial reuse. Unlike conventional synchronization approaches that assume static point to point links, the proposed framework explicitly captures the coupling between spatial beam geometry, photon reception statistics, and synchronization accuracy. Particularly, we developed a comprehensive system model that integrates photon pair generation, grid based beam delivery, spatial uncertainty, and realistic detector behaviour. The analysis showed that synchronization performance is fundamentally determined by the number of successfully detected photon pairs, which is shaped by beam alignment and user position uncertainty. Through a Cramér Rao based formulation, we also established a direct quantitative relationship between spatial design parameters and achievable timing precision.
The impact of practical impairments was further examined through detailed modeling and simulation. Multipath propagation was found to introduce a bounded timing bias rather than instability, while non Gaussian detector jitter primarily affects convergence speed rather than steady state accuracy. These effects define fundamental performance limits but do not undermine the viability of the proposed approach. Even under realistic indoor channel conditions, the framework achieves synchronization accuracy on the order of a few picoseconds. Furthermore, the computational feasibility of the method was demonstrated. The two stage synchronization algorithm supports efficient FFT based implementation with modest memory requirements. Thus, the prposed algorithm is suitable for real time operation.
Several research directions can be built upon this work. For instance, future efforts may include experimental validation using a hardware testbed, adaptive calibration techniques to mitigate multipath induced bias, and extensions to multi user and multi cell scenarios. In addition, integration with hybrid RF optical architectures and higher layer network protocols can be explored to enable fully deployable quantum enabled indoor communication systems.